# MMHMR: Generative Masked Modeling for Hand Mesh Recovery

### v1

@@ Appendix I Impact of Pose Tokenizer on MMHMR
Appendix J Impact of Multi-Scale Features
The ablation study on multi-scale feature resolutions in MMHMR (as shown in Table 15) highlights the trade-off between accuracy and computational cost. Including resolutions up to 4×4$\times$4 × yields slight accuracy gains, with PA-MPJPE reducing to 7.0 mm on HO3Dv3 and 5.7 mm on FreiHAND. However, the addition of higher resolutions, such as 1×1$\times$1 × and 2×2$\times$2 ×, results in inconsistent or degraded performance. Specifically, the inclusion of 1×,4×,8×1$\times$,4$\times$,8$\times$1 × , 4 × , 8 × scales increases PA-MPJPE to 7.2 mm on HO3Dv3 and 5.8 mm on FreiHAND. Adding 2×2$\times$2 × further worsens performance, reaching 7.6 mm on HO3Dv3 and 6.1 mm on FreiHAND, while significantly increasing computational overhead. Notably, the omission of lower-scale features (e.g., 1×1$\times$1 ×) leads to performance degradation, highlighting the importance of combining fine-grained details with holistic structure. While multi-scale features remain critical, the study demonstrates that not all resolutions contribute equally, with 1×1$\times$1 × and 4×4$\times$4 × emerging as the optimal balance for accuracy and computational efficiency.
@@
Appendix K Cross-Attention Layers in MMHMR
The ablation study in Table 16 highlights the pivotal role of Deformable Cross-Attention Layers in the Context-Infused Masked Synthesizer of MMHMR. Increasing layers from 2 to 4 yields significant performance gains, reducing PA-MPJPE to 5.7 mm and PA-MPVPE to 5.5 mm on the FreiHAND dataset. This improvement underscores the layers’ effectiveness in fusing multi-scale contextual features and refining token dependencies for enhanced 3D hand mesh reconstruction. However, further increasing the layers beyond 4 results in diminishing returns, with slight performance degradation at 6 and 8 layers (e.g., PA-MPVPE increases to 5.8 mm and 6.1 mm, respectively). This decline suggests that additional layers introduce unnecessary complexity, potentially overfitting or disrupting the model’s ability to generalize effectively. The findings reveal that 4 layers provide the optimal balance, leveraging the benefits of cross-attention mechanisms without incurring computational overhead or accuracy trade-offs.
@@
Appendix L Qualitative Results
@@
Comparison of State-of-the-Art (SOTA) Methods. Figure 12 demonstrates the superiority of MMHMR over other SOTA methods in recovering 3D hand meshes. Unlike competing approaches, MMHMR employs a generative masked modeling framework, enabling it to synthesize unobserved or occluded hand regions. This capability allows MMHMR to achieve robust and precise 3D reconstructions, even in scenarios with heavy occlusions, intricate hand-object interactions, or diverse hand poses. By refining masked tokens, MMHMR effectively addresses ambiguities in the 2D-to-3D mapping process, resulting in highly accurate reconstructions.
MMHMR’s Performance on In-the-Wild Images. Figure 13 highlights MMHMR’s robustness in real-world conditions. The model demonstrates its ability to recover accurate 3D hand meshes from single RGB images, excelling in challenging scenarios such as occlusions, hand-object interactions, and diverse hand appearances. This versatility underscores MMHMR’s applicability to real-world tasks, where robust and reliable performance is essential.
Challenging Poses from the HInt Benchmark. Figure 14 illustrates MMHMR’s effectiveness in reconstructing 3D hand meshes for challenging poses from the HInt Benchmark [46]. The model accurately handles extreme articulations and unconventional hand configurations, showcasing its ability to generalize to complex datasets and produce high-fidelity results.
Reference Key Points in the Deformable Cross-Attention. Figure 15 visualizes the interaction between reference keypoints (yellow) and sampling offsets (red) in the Deformable Cross-Attention module of MMHMR’s Masked Synthesizer. By leveraging 2D pose as guidance, the model dynamically samples and refines critical features for accurate 3D reconstruction. This mechanism proves crucial in handling severe occlusions, intricate hand-object interactions, and complex viewpoints, ensuring precise alignment between 2D observations and 3D predictions.
Impact of Pose Token Confidence on MMHMR Reconstruction. Figure 16 explores the impact of pose token confidence on reconstruction accuracy. The comparison of 3D hand meshes generated using different levels of token confidence reveals that high-confidence tokens (Top-1) produce reconstructions closely aligned with the ground truth. In contrast, lower-confidence tokens introduce distortions, highlighting the importance of prioritizing high-confidence tokens for achieving accurate and robust 3D hand mesh reconstructions.
@@
Figure 12: Comparison of State-of-the-Art (SOTA) methods for 3D hand mesh recovery, highlighting the performance of MMHMR. Unlike other approaches, MMHMR employs a generative masked modeling framework to synthesize unobserved or occluded regions, enabling precise and robust 3D hand reconstructions even in challenging scenarios such as heavy occlusions, hand-object interactions, and diverse hand poses. This comparison underscores MMHMR’s ability to outperform competing methods by addressing ambiguities in the 2D-to-3D mapping process through its innovative masked token refinement strategy.
@@
Figure 13: MMHMR’s performance on in-the-wild images, demonstrating its ability to recover accurate and robust 3D hand meshes from single RGB inputs. The model excels in challenging scenarios, including occlusions, hand-object interactions, and diverse hand appearances, showcasing its versatility and reliability in real-world conditions.
@@
Figure 15: Visualization of reference key points (yellow) and sampling offsets (red) in the Deformable Cross-Attention module of MMHMR’s Masked Synthesizer. The 2D pose acts as a guidance signal, enabling the model to dynamically sample and refine features critical for reconstructing accurate 3D hand meshes (rightmost column). This mechanism adapts to challenging scenarios, such as severe occlusions, intricate hand-object interactions, and complex viewpoints, ensuring precise alignment between 2D observations and 3D predictions for robust and high-fidelity reconstructions.
Figure 16: Impact of pose token confidence on MMHMR’s reconstruction accuracy. The figure compares 3D hand meshes using the 1st, 7th, 10th, 100th, 250th, and 1000th most confident tokens. High-confidence tokens (Top-1) yield reconstructions closely aligned with the ground truth, while lower-confidence tokens introduce distortions, underscoring the importance of prioritizing high-confidence tokens for accurate 3D hand mesh reconstruction.

### v2

@@ Appendix I Masking Ratio during Training
Appendix M Impact of Multi-Scale Features
The ablation study on multi-scale feature resolutions in MaskHand (as shown in Table 16) highlights the trade-off between accuracy and computational cost. Including resolutions up to 4×4$\times$4 × yields slight accuracy gains, with PA-MPJPE reducing to 7.0 mm on HO3Dv3 and 5.7 mm on FreiHAND. However, the addition of higher resolutions, such as 1×1$\times$1 × and 2×2$\times$2 ×, results in inconsistent or degraded performance. Specifically, the inclusion of 1×,4×,8×1$\times$,4$\times$,8$\times$1 × , 4 × , 8 × scales increases PA-MPJPE to 7.2 mm on HO3Dv3 and 5.8 mm on FreiHAND. Adding 2×2$\times$2 × further worsens performance, reaching 7.6 mm on HO3Dv3 and 6.1 mm on FreiHAND, while significantly increasing computational overhead. Notably, the omission of lower-scale features (e.g., 1×1$\times$1 ×) leads to performance degradation, highlighting the importance of combining fine-grained details with holistic structure. While multi-scale features remain critical, the study demonstrates that not all resolutions contribute equally, with 1×1$\times$1 × and 4×4$\times$4 × emerging as the optimal balance for accuracy and computational efficiency.
@@
Appendix N Deformable Cross-Attention Layers in MaskHand
The ablation study in Table 17 highlights the pivotal role of Deformable Cross-Attention Layers in the Context-Infused Masked Synthesizer of MaskHand. Increasing layers from 2 to 4 yields significant performance gains, reducing PA-MPJPE to 5.7 mm and PA-MPVPE to 5.5 mm on the FreiHAND dataset. This improvement underscores the layers’ effectiveness in fusing multi-scale contextual features and refining token dependencies for enhanced 3D hand mesh reconstruction. However, further increasing the layers beyond 4 results in diminishing returns, with slight performance degradation at 6 and 8 layers (e.g., PA-MPVPE increases to 5.8 mm and 6.1 mm, respectively). This decline suggests that additional layers introduce unnecessary complexity, potentially overfitting or disrupting the model’s ability to generalize effectively. The findings reveal that 4 layers provide the optimal balance, leveraging the benefits of cross-attention mechanisms without incurring computational overhead or accuracy trade-offs.
@@
Appendix O Qualitative Results in the Wild
@@
Comparison of State-of-the-Art (SOTA) Methods. Figure 13 demonstrates the superiority of MaskHand over other SOTA methods in recovering 3D hand meshes. Unlike competing approaches, MaskHand employs a generative masked modeling framework, enabling it to synthesize unobserved or occluded hand regions. This capability allows MaskHand to achieve robust and precise 3D reconstructions, even in scenarios with heavy occlusions, intricate hand-object interactions, or diverse hand poses. By refining masked tokens, MaskHand effectively addresses ambiguities in the 2D-to-3D mapping process, resulting in highly accurate reconstructions.
@@
Figure 13: Comparison of State-of-the-Art (SOTA) methods for 3D hand mesh recovery, highlighting the performance of MaskHand. Unlike other approaches, MaskHand employs a generative masked modeling framework to synthesize unobserved or occluded regions, enabling precise and robust 3D hand reconstructions even in challenging scenarios such as heavy occlusions, hand-object interactions, and diverse hand poses. This comparison underscores MaskHand’s ability to outperform competing methods by addressing ambiguities in the 2D-to-3D mapping process through its innovative masked token refinement strategy.
Multiple Reconstruction Hypotheses with Explicit Confidence Levels. Figure  14 and  15 illustrates MaskHand’s 3D hand mesh reconstructions in occluded scenarios, ranked by confidence. The comparison of reconstructions across different confidence levels reveals that high-confidence hypotheses produce meshes that closely align with the ground truth, ensuring structural accuracy and fidelity. As confidence decreases (e.g., from the 100th to the 1000th hypothesis), the reconstructions degrade, exhibiting distortions and unrealistic poses. This highlights the significance of MaskHand’s confidence-aware modeling, where prioritizing high-confidence hypotheses leads to more accurate and robust 3D hand reconstructions.
Figure 14: Multiple reconstruction hypotheses with explicit confidence levels. The figure illustrates MaskHand’s 3D hand mesh reconstructions in occluded scenarios, ranked by confidence. High-confidence hypotheses closely align with the ground truth, ensuring structural accuracy and fidelity. As confidence decreases (e.g., from the 100th to the 1000th hypothesis), reconstructions degrade, exhibiting distortions and unrealistic poses. This highlights the importance of prioritizing high-confidence hypotheses for robust and accurate 3D hand reconstruction.
Figure 15: Multiple reconstruction hypotheses with explicit confidence levels. The figure illustrates MaskHand’s 3D hand mesh reconstructions in occluded scenarios, ranked by confidence. High-confidence hypotheses closely align with the ground truth, ensuring structural accuracy and fidelity. As confidence decreases (e.g., from the 100th to the 1000th hypothesis), reconstructions degrade, exhibiting distortions and unrealistic poses. This highlights the importance of prioritizing high-confidence hypotheses for robust and accurate 3D hand reconstruction.
Reference Key Points in the Deformable Cross-Attention. Figure 16 visualizes the interaction between reference keypoints (yellow) and sampling offsets (red) in the Deformable Cross-Attention module of MaskHand’s Masked Synthesizer. By leveraging 2D pose as guidance, the model dynamically samples and refines critical features for accurate 3D reconstruction. This mechanism proves crucial in handling severe occlusions, intricate hand-object interactions, and complex viewpoints, ensuring precise alignment between 2D observations and 3D predictions.
Figure 16: Visualization of reference key points (yellow) and sampling offsets (red) in the Deformable Cross-Attention module of MaskHand’s Masked Synthesizer. The 2D pose acts as a guidance signal, enabling the model to dynamically sample and refine features critical for reconstructing accurate 3D hand meshes (rightmost column). This mechanism adapts to challenging scenarios, such as severe occlusions, intricate hand-object interactions, and complex viewpoints, ensuring precise alignment between 2D observations and 3D predictions for robust and high-fidelity reconstructions.
@@
MaskHand’s Performance on In-the-Wild Images. Figure 17 highlights MaskHand’s robustness in real-world conditions. The model demonstrates its ability to recover accurate 3D hand meshes from single RGB images, excelling in challenging scenarios such as occlusions, hand-object interactions, and diverse hand appearances. This versatility underscores MaskHand’s applicability to real-world tasks, where robust and reliable performance is essential.
Figure 17: MaskHand’s performance on in-the-wild images, demonstrating its ability to recover accurate and robust 3D hand meshes from single RGB inputs. The model excels in challenging scenarios, including occlusions, hand-object interactions, and diverse hand appearances, showcasing its versatility and reliability in real-world conditions.
Challenging Poses from the HInt Benchmark. Figure  18 and  19 illustrates MaskHand’s effectiveness in reconstructing 3D hand meshes for challenging poses from the HInt Benchmark [49]. The model accurately handles extreme articulations and unconventional hand configurations, showcasing its ability to generalize to complex datasets and produce high-fidelity results.